Using annihilator method obtain particular solution of y"'+y= cosx+sinx


1
Expert's answer
2021-08-19T14:04:02-0400

Corresponding homogeneous differential equation


"y'''+y=0"

Characteristic equarion


"r^3+1=0"

"(r+1)(r^2-r+1)=0"

"r_1=-1, r_2=\dfrac{1}{2}-\dfrac{\sqrt{3}}{2}i, r_3=\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i"

The general solution of the homogeneous differential equation


"y_h=c_1e^{-x}+c_2e^{x/2}\sin (\dfrac{\sqrt{3}}{2}x)+c_3e^{x/2}\cos (\dfrac{\sqrt{3}}{2}x)"




"(D^3+1)y=\cos x+\sin x"

Find the particular solution of the nonhomogeneous differential equation


"y_p=A\cos x+B\sin x"

Then


"y_p'=-A\sin x+B\cos x"

"y_p''=-A\cos x-B\sin x"




"y_p'''=A\sin x-B\cos x"

Substitute


"A\sin x-B\cos x+A\cos x+B\sin x="

"=\cos x+\sin x"


"A+B=1"

"-B+A=1"

"A=1"

"B=0"

The particular solution of the nonhomogeneous differential equation is


"y_p=\cos x"